# Navigating Extremes: Dynamic Sparsity in Large Output Spaces

### v2

@@ 1 Introduction
Recent research Frankle and Carbin (2018); Frankle et al. (2020); Malach et al. (2020); Chen et al. (2021a) has demonstrated that densely-connected neural networks contain sparse subnetworks — often dubbed “winning lottery tickets” — that can deliver performance comparable to the full networks but with substantially reduced compute and memory demands. Unlike conventional techniques that start with a trained dense model and employ iterative pruning or one-shot pruning, Dynamic Sparse Training (DST) Mocanu et al. (2018); Evci et al. (2020); Liu et al. (2021a) initializes the a sparse architecture and dynamically explores subnetwork configurations through periodic pruning and regrowth, typically informed with heuristic saliency criteria such as weight and gradient magnitudes. This approach is particularly advantageous in scenarios constrained by a fixed memory budget during the training phase, making DST viable across various domains Chen et al. (2021a); Dietrich et al. (2021); Liu et al. (2021b). For instance, in reinforcement learning Tan et al. (2022); Graesser et al. (2022), DST has been shown to significantly outperform traditional dense models. Additionally, models trained using DST often exhibit enhanced robustness Guo et al. (2018); Özdenizci and Legenstein (2021); Chen et al. (2021b); Diffenderfer et al. (2021); Liu et al. (2022). However, the application of DST comes with challenges, notably prolonged training times; for example, RigL Evci et al. (2020) and ITOP Liu et al. (2021a) require up to five and two times as many optimization steps during training, respectively, to match the generalisation performance of dense networks at high sparsity levels ($\geq 80\%$). The prolonged training time in these works is often linked to the need for in-time overparameterization Liu et al. (2021a) and poor gradient flow in sparse networks. Recent advances Evci et al. (2022); Price and Tanner (2021); Curci et al. (2021) aimed at improving gradient flow have been introduced to mitigate these extended training durations, enhancing the practicality of DST methodologies.
@@
In fact, direct application of existing DST methods yields unsatisfactory results on XMC tasks due to typically noisy data and poor gradient signal propagation through the sparse classifier, slowing training convergence to an extent that it is not practically useful. Consequently, we follow Schultheis and Babbar (2023) and adapt the model architecture by integrating an intermediate layer that is larger than the embedding from the encoder but still significantly smaller than the final output layer. While this was found to be sufficient to achieve good results with fixed encodings, it fails if the encoder is a trainable transformer (Devlin et al., 2018; Sanh, 2019) for label spaces with more than one hundred thousand elements, particularly at high levels of sparsity. The primary challenge arises from the noisy gradients prevalent at the onset of training, which are inadequate for guiding the fine-tuning of the encoder effectively. To mitigate this issue, we introduce an auxiliary loss. This loss uses a more coarse-grained objective, assigning instances to clusters of labels, where scores for each cluster are calculated using a dense classification layer. This auxiliary component stabilizes the gradient flow and enhances the encoder’s adaptability during the critical early phases of training and is turned off during later epochs to not interfere with the main task. Figure 1 illustrates the architectural changes that ensure good training performance at different label space sizes and sparsity levels.
@@
  • Enhancements in training efficiency: We propose novel modifications to the conventional DST framework that significantly curtail training durations while concurrently delivering competitive performance metrics when benchmarked against dense model baselines and other specialized XMC methodologies. These enhancements are pivotal in demonstrating DST’s scalability and efficiency to large label spaces.
@@ 2.1 Background
The XMC models are typically comprised of two main components: (i) an encoder $\mathcal{E}_{\theta}:\chi\rightarrow\mathbb{R}^{d}$, which embeds data points into a $d$-dimensional real space, primarily utilizing a transformer architecture (Vaswani et al., 2017) and (ii) A One-vs-All classifier $W=\{w_{l}\}_{l\in[L]}$, where $w_{l}$ denotes the classifier for label $l$, integrated as the last layer of the neural network in end-to-end training settings. In a typical DST scenario, one would sparsify the language model used as the encoder, potentially even leaving the classifier fully dense (Kurtic et al., 2024). However, in XMC, most of the networks weights are in the classifier layer, so in order to achieve a reduction in memory consumption, its weight matrix $W_{s}=\{w_{l}^{s}\}_{l\in[L]}$ must be sparsified.
This sparse layer $W_{s}$ is then periodically updated in a prune-regrow-train loop, that is, every $\Delta T$ steps, a fraction of active weights are pruned and the same number of inactive weights are regrown. The updated sparse topology is then trained with regular gradient descent for the next $\Delta T$ steps. There are many possible choices for pruning and regrowth criteria (Hoefler et al., 2021); to keep memory consumption low, however, we need to choose a method that does not require auxiliary buffers proportional to the size of the dense layer. This excludes methods such as requiring second-order information (Hassibi and Stork, 1992), or tracking of dense gradients or other per-weight information (Dai et al., 2019; Dettmers and Zettlemoyer, 2019; Lin et al., 2020). Evci et al. (2020) argue that RigL only needs dense gradients in an ephemeral capacity — they can be discarded as soon as the regrowth step for the current layer is done, but before the regrow step of the next layer is started — but in the XMC setup, the prohibitively large memory consumption arises already from a single layer. Therefore, we select magnitude-based pruning and random regrowth (Mocanu et al., 2018). Magnitude-based pruning has been shown to be a remarkably strong baseline (Nowak et al., 2024).
@@ 2.2 Memory-Efficient Training: Fixed Fan-In Sparse Layer
Unstructured sparsity is notoriously difficult to speed-up on GPUs Gale et al. (2020), and consequently most DST studies simulate sparsity by means of a binary mask (Curci et al., 2021; Lee et al., 2023). On the other hand, highly structured sparsity, such as 2:4 sparsity (Castro et al., 2023), enjoys hardware acceleration and memory reduction (Mishra et al., 2021), but results in deteriorated model accuracy compared to unstructured sparsity (Liu and Wang, 2023). As a compromise, semi-structured sparsity (Lasby et al., 2024; Schultheis and Babbar, 2023) imposes a fixed fan-in to each neuron. This eliminates work imbalances between different neurons, leading to an efficient and simple storage format for sparse weights, where each sparse weight needs only a single integer index. For 32-bit floating point weights with 16-bit indices (i.e., at most 65k features in the embedding layer), this leads to a 50% storage overhead for sparse weights; however, for training, gradient and two momentum terms are needed, which share the same indexing structure, reducing the effective overhead to just 12.5%.
@@ 3.3 Empirical performance
Table 2 presents our primary results on the datasets, compared with the aforementioned baselines. The performance metrics for XMC baselines are reported from their original papers. However, for peak memory consumption, we re-ran these baselines in half precision with the same batch size, as all baselines are also evaluated in half precision. Following DST protocols, we extended the training duration for RigL, Dense Bottleneck, and our method to twice the number of steps used for dense models. Certain baselines (referred to as OOM) could not scale to the Amazon-3M dataset. Our results demonstrate that our method significantly reduces memory usage while maintaining competitive performance. On the Amazon-3M dataset, our approach delivers comparable performance to dense models while achieving a 3.4-fold reduction in memory usage and a 5.-fold reduction compared to the XMC baseline. Furthermore, within the memory-efficient model regime, our method consistently outperforms the Dense Bottleneck model. To further validate the robustness of our approach, we evaluated it on the label features datasets, as shown in Table 3. Notably, as the label space size increases, we need to adjust to a comparatively lower sparsity to maintain performance, discussed in detail in subsequent sections.
@@ Appendix B Baselines and Related Works
  • Structured DST Lasby et al. (2024); Schultheis and Babbar (2023): In contrast to RigL, structured DST methods adds constraints to the learned network topology such that the network is more amenable to acceleration on commodity GPUs. In our case, we employ the fixed fan-in constraint which reduces both the latency and memory consumption of the final classification layer in the XMC setting. Further, since the dense gradient information is not available in the XMC task, we simply randomly regrow weights as per to SET (Mocanu et al., 2018) which has proven to be a robust baseline in the DST literature.

### v3

@@ 1 Introduction
Recent research Frankle and Carbin (2018); Frankle et al. (2020); Malach et al. (2020); Chen et al. (2021a) has demonstrated that densely-connected neural networks contain sparse subnetworks — often dubbed “winning lottery tickets” — that can deliver performance comparable to the full networks but with substantially reduced compute and memory demands. Unlike conventional techniques that start with a trained dense model and employ iterative pruning or one-shot pruning, Dynamic Sparse Training (DST) Mocanu et al. (2018); Evci et al. (2020); Liu et al. (2021a) initializes a sparse architecture and dynamically explores subnetwork configurations through periodic pruning and regrowth, typically informed with heuristic saliency criteria such as weight and gradient magnitudes. This approach is particularly advantageous in scenarios constrained by a fixed memory budget during the training phase, making DST viable across various domains Chen et al. (2021a); Dietrich et al. (2021); Liu et al. (2021b). For instance, in reinforcement learning Tan et al. (2022); Graesser et al. (2022), DST has been shown to significantly outperform traditional dense models. Additionally, models trained using DST often exhibit enhanced robustness Guo et al. (2018); Özdenizci and Legenstein (2021); Chen et al. (2021b); Diffenderfer et al. (2021); Liu et al. (2022). However, the application of DST comes with challenges, notably prolonged training times; for example, RigL Evci et al. (2020) and ITOP Liu et al. (2021a) require up to five and two times as many optimization steps during training, respectively, to match the generalisation performance of dense networks at high sparsity levels ($\geq 80\%$). The prolonged training time in these works is often linked to the need for in-time overparameterization Liu et al. (2021a) and poor gradient flow in sparse networks. Recent advances Evci et al. (2022); Price and Tanner (2021); Curci et al. (2021); Rajeshkumar Bambhaniya et al. (2024) aimed at improving gradient flow have been introduced to mitigate these extended training durations, enhancing the practicality of DST methodologies.
@@
In fact, direct application of existing DST methods yields unsatisfactory results on XMC tasks due to typically noisy data and poor gradient signal propagation through the sparse classifier, slowing training convergence to an extent that it is not practically useful. Consequently, we follow Schultheis and Babbar (2023) and adapt the model architecture by integrating an intermediate layer that is larger than the embedding from the encoder but still significantly smaller than the final output layer. While this was found to be sufficient to achieve good results with fixed encodings, if the encoder is a trainable transformer (Devlin et al., 2018; Sanh, 2019) for label spaces with more than one hundred thousand elements, particularly at high levels of sparsity. The primary challenge arises from the noisy gradients prevalent at the onset of training, which are inadequate for guiding the fine-tuning of the encoder effectively. To mitigate this issue, we introduce an auxiliary loss. This loss uses a more coarse-grained objective, assigning instances to clusters of labels, where scores for each cluster are calculated using a dense classification layer. This auxiliary component stabilizes the gradient flow and enhances the encoder’s adaptability during the critical early phases of training and is turned off during later epochs to not interfere with the main task. Figure 1 illustrates the architectural changes that ensure good training performance at different label space sizes sparsity levels.
@@
  • Enhancements in training efficiency: We propose novel modifications to the conventional DST framework that significantly curtail training durations while delivering competitive performance metrics when benchmarked against dense model baselines and other specialized XMC methodologies. These enhancements are pivotal in demonstrating DST’s scalability and efficiency to large label spaces.
@@ 2.1 Background
XMC models are typically comprised of two main components: (i) an encoder $\mathcal{E}_{\theta}:\chi\rightarrow\mathbb{R}^{d}$, which embeds data points into a $d$-dimensional real space, primarily utilizing a transformer architecture (Vaswani et al., 2017) and (ii) A One-vs-All classifier $W=\{w_{l}\}_{l\in[L]}$, where $w_{l}$ denotes the classifier for label $l$, integrated as the last layer of the neural network in end-to-end training settings. In a typical DST scenario, one would sparsify the language model used as the encoder, potentially even leaving the classifier fully dense (Kurtic et al., 2024). However, in XMC, most of the networks weights are in the classifier layer, so in order to achieve a reduction in memory consumption, its weight matrix $W_{s}=\{w_{l}^{s}\}_{l\in[L]}$ must be sparsified.
This sparse layer $W_{s}$ is then periodically updated in a prune-regrow-train loop, that is, every $\Delta T$ steps, a fraction of active weights is pruned and the same number of inactive weights are regrown. The updated sparse topology is then trained with regular gradient descent for the next $\Delta T$ steps. There are many possible choices for pruning and regrowth criteria (Hoefler et al., 2021); to keep memory consumption low, however, we need to choose a method that does not require auxiliary buffers proportional to the size of the dense layer. This excludes methods such as requiring second-order information (Hassibi and Stork, 1992), or tracking of dense gradients or other per-weight information (Dai et al., 2019; Dettmers and Zettlemoyer, 2019; Lin et al., 2020). Evci et al. (2020) argue that RigL only needs dense gradients in an ephemeral capacity — they can be discarded as soon as the regrowth step for the current layer is done, but before the regrow step of the next layer is started — but in the XMC setup, the prohibitively large memory consumption arises already from a single layer. Therefore, we select magnitude-based pruning and random regrowth (Mocanu et al., 2018). Magnitude-based pruning has been shown to be a remarkably strong baseline (Nowak et al., 2024).
@@ 2.2 Memory-Efficient Training: Fixed Fan-In Sparse Layer
Unstructured sparsity is notoriously difficult to speed-up on GPUs Gale et al. (2020), and consequently most DST studies simulate sparsity by means of a binary mask (Curci et al., 2021; Lee et al., 2023). On the other hand, highly structured sparsity, such as 2:4 sparsity (Castro et al., 2023), enjoys hardware acceleration and memory reduction (Mishra et al., 2021), but may result in deteriorated model accuracy compared to unstructured sparsity (Liu and Wang, 2023). As a compromise, semi-structured sparsity (Muralidharan, 2023; Lasby et al., 2024; Schultheis and Babbar, 2023) imposes a fixed fan-in to each neuron. This eliminates work imbalances between different neurons, leading to an efficient and simple storage format for sparse weights, where each sparse weight needs only a single integer index, resulting in ELLPACK format (Rice and Boisvert, 1985) without any padding. For 32-bit floating point weights with 16-bit indices (i.e., at most 65k features in the embedding layer), this leads to a 50% storage overhead for sparse weights; however, for training, gradient and two momentum terms are needed, which share the same indexing structure, reducing the effective overhead to just 12.5%.
@@ 3.3 Empirical performance
Table 2 presents our primary results on the datasets, compared with the aforementioned baselines. The performance metrics for XMC baselines are reported from their original papers. However, for peak memory consumption, we re-ran these baselines in half precision with the same batch size, as all baselines are also evaluated in half precision. Following DST protocols, we extended the training duration for RigL, Dense Bottleneck, and our method to twice the number of steps used for dense models. Certain baselines (referred to as OOM) could not scale to the Amazon-3M dataset. Our results demonstrate that our method significantly reduces memory usage while maintaining competitive performance. On the Amazon-3M dataset, our approach delivers comparable performance to dense models while achieving a 3.4-fold reduction in memory usage and a 5-fold reduction compared to the XMC baseline. Furthermore, within the memory-efficient model regime, our method consistently outperforms the Dense Bottleneck model. To further validate the robustness of our approach, we evaluated it on the label features datasets, as shown in Table 3. Notably, as the label space size increases, we need to adjust to a comparatively lower sparsity to maintain performance, discussed in detail in subsequent sections.
@@ Appendix B Baselines and Related Works
  • Structured DST Muralidharan (2023); Lasby et al. (2024); Schultheis and Babbar (2023): In contrast to RigL, structured DST methods adds constraints to the learned network topology such that the network is more amenable to acceleration on commodity GPUs. In our case, we employ the fixed fan-in constraint which reduces both the latency and memory consumption of the final classification layer in the XMC setting. Further, since the dense gradient information is not available in the XMC task, we simply randomly regrow weights as per to SET (Mocanu et al., 2018) which has proven to be a robust baseline in the DST literature.
@@ Appendix F Computational Resources
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and precede the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction provide a suffcient explanation of the main idea of the paper. The problem setting and the contributions are explicitly stated in the introduction section of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: A section discussing the limitations has been added towards the end of the paper. Various computational considerations are the main part of the paper, and are adequately discussed.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: There are no formal results proved or claimed in this paper
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The dataset details, architecture outline and hyper-parameter details are sufficiently explained in the main body of the paper and appendices.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: Data-[Yes] , and code - [No]
  24. Justification: We perform experiments on publicly available datasets. The code will be made publicly available in the near future.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental setting and hyperparameter details are sufficiently explained in the main paper and appendices.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: While the train, and test splits are standard, due to the scale of datasets, and computations involved performing significance tests is not undertaken by the research community in this domain.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: This has been sufficiently explained in the experiments section of the paper.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Ethical considerations
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: A brief description of the societal impact of work in provided in the paper.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not release any Language model, or image generators.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The link to the repository hosting publicly available datasets has been provided.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release any new assests.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No study involving humans was part of this work.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No study involving humans was part of this work.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.